1/ The following information ($ in millions) comes from a recent annual report of Amazon.com, Inc.:

Net sales $ 10,723
Total assets 4,427
End of year balance in cash 1,218
Total stockholders' equity 525
Gross profit (Sales – Cost of Sales) 2,649
Net increase in cash for the year 10
Operating expenses 2,067
Net operating cash flow 801
Other income (expense), net (21)

  

Compute the income before income tax for Amazon.

Multiple Choice

  • $ 481

  • $ 681

  • $ 561

  • $ 641

  • 2/ The following information ($ in millions) comes from a recent annual report of Amazon.com, Inc.:

    Net sales $ 10,786
    Total assets 4,494
    End of year balance in cash 1,130
    Total stockholders' equity 420
    Gross profit (Sales – Cost of Sales) 2,519
    Net increase in cash for the year 20
    Operating expenses 2,063
    Net operating cash flow 603
    Other income (expense), net (16)

    Compute Amazon's cost of goods sold for the year.

    Multiple Choice

  • $ 7,160

  • $ 4,471

  • $ 5,097

  • $ 8,267

  • 3/ Alpaca Corporation had revenues of $270,000 in its first year of operations. The company has not collected on $19,600 of its sales and still owes $26,000 on $80,000 of merchandise it purchased. The company had no inventory on hand at the end of the year. The company paid $13,000 in salaries. Owners invested $29,000 in the business and $29,000 was borrowed on a five-year note. The company paid $3,100 in interest that was the amount owed for the year, and paid $7,000 for a two-year insurance policy on the first day of business. Alpaca has an effective income tax rate of 40%.


    Compute net income for the first year for Alpaca Corporation.

    Multiple Choice

  • $ 113,140

  • $ 102,240

  • $ 170,400

  • $ 190,000

  • 4/ Land was acquired in 2018 for a future building site at a cost of $40,700. The assessed valuation for tax purposes is $28,100, a qualified appraiser placed its value at $49,800, and a recent firm offer for the land was for a cash payment of $44,300. The land should be reported in the financial statements at:

    Multiple Choice

  • $ 28,100.

  • $ 40,700.

  • $ 44,300.

  • $ 49,800.
Answer #1

1) Income before income tax = 2649-2067-21 = 561

2) Cost of goods sold = 10786-2519 = 8267

3)

Calculation of net income for the first year for Alpaca Corporation
Revenues 270,000
Less: Cost of Merchandise used (80,000)
Less: Salaries (13,000)
Less: Interest on Notes (3,100)
Less: Insurance Expense = 7000 /2 (3,500)
Income before tax 170,400
Less: Tax @40% (68,160)
Income after tax 102,240

4) The land should be reported in the financial statements at the price at which it was acquired $40,700.